OpenStudy (anonymous):

Can someone help me? I have 5 problems and they all have to do with fractions...

OpenStudy (anonymous):

\[7\frac{ 3 }{ 4 } x \frac{ 2 }{ 3 } - \frac{ 5 }{ 12 }\]